OpenStudy (anonymous):

Solve Write your solution using set builder notation. Please show all of your work -8-11(3x+7)<40x-9

OpenStudy (anonymous):

\[\le\] is suppose to be in the problem

OpenStudy (aroub):

-8-33x-77<40x-9 -40x-33x<77+8-9 -73x<76 x> -76/73 \[\S.\S= {x \in \mathbb{R}| x \ge -76/73} \]